### v1

# Learning Noise-Robust Stable Koopman Operator for Control with Physics-Informed Observables

Abstract
We propose a novel learning framework for Koopman operator of nonlinear dynamical systems that is informed by the governing equation and guarantees long-time stability and robustness to noise. In contrast to existing frameworks where either ad-hoc observables or blackbox neural networks are used to construct observables in the extended dynamic mode decomposition (EDMD), our observables are informed by governing equations via Polyflow. To improve the noise robustness and guarantee long-term stability, we designed a stable parameterization of the Koopman operator together with a progressive learning strategy for roll-out recurrent loss. To further improve model performance in the phase space, a simple iterative strategy of data augmentation was developed. Numerical experiments of prediction and control of classic nonlinear systems with ablation study showed the effectiveness of the proposed techniques over several state-of-the-art practices.
I Introduction
Data-driven methods have shown promising potential in modeling and control of nonlinear systems [1, 2]. Interested readers are encouraged to read several excellent reviews of Koopman operator [1, 3, 4, 5, 6]. The Koopman operator has been applied across diverse fields, including robust COVID-19 prediction [7], reinforcement learning [8], soft robotics control [9], stability analysis of genetic toggle switches [10], and modeling turbulent shear flows [11].
The biggest benefit of Koopman operator for dynamics and control community is to facilitate the use of well-established linear control/observer techniques for the nonlinear systems[12, 13, 14], provided that a good approximation of global linear representation from data is learned. For example, model predictive control (MPC) is typically computationally expensive due to the need to solve non-convex optimization online. Traditional linearization techniques that approximate the nonlinear system around a local reference point might result in suboptimal MPC performance and could struggle with significant deviations from this point. On the other hand, the linear representation from Koopman operator can be used to transform the non-convex optimization of the nonlinear MPC into convex optimization of a linear MPC as shown by [15]. In addition to classical MPC, such linear representation can also help LQR [13], anomaly detection [16] and fault tolerant control [17] to nonlinear systems.
Despite the several successful application of Koopman operator, there still remain several challenges. Colbrook et al. [18] examine the limitations of the Koopman operator and systematically determines when and how Koopman spectral properties can be learned. Wu et al. [19] reveal that under mild noise and nonlinearity, DMD can struggle to accurately recover and predict dynamics, even with an exact linear model.
Among various challenges, the first challenge is finding an appropriate invariant subspace for the Koopman operator. This involves selecting suitable basis functions that span an invariant subspace. Common choices of basis functions include radial basis functions (RBFs) [20] or polynomial functions [21]. Another approach is to learn the observables from dictionary of functions [22], or identifying dominant eigenfunctions from data [13]. Otto et al. [23] formulate the dynamics of observables under the Koopman generator as a bilinear hidden Markov model. They determine the model parameters using the expectation-maximization (EM) algorithm, thereby eliminating the need to explicitly specify a dictionary of functions, as required by existing methods like Extended Dynamic Mode Decomposition (EDMD). However, in order to lift the system, one has to perform a non-trivial joint optimization of hidden state and dynamics.
Recently, the concept of Polyflow was introduced by [24]. Polyflow employs the governing equations of the system to construct basis functions, offering a non-local linearization approach for nonlinear differential equations. This concept, which we refer to as physics-informed observables, is closely related to the derivative-based approximation of Koopman operator [25] except closure is not ignored in Polyflow. Despite the success of initial attempt of Polyflow for MPC [26], simply minimizing the one-step ahead error will be prone to noisy data.
The second challenge in learning Koopman operator from data is ensuring long-term prediction accuracy. The use of roll-out techniques for recurrent loss has proven beneficial when data is limited and noisy both in the deep learning community [27, 28, 29] and the reduced-order modeling community [30]. However, optimizing the recurrent loss can be tricky and prone to poor local minimum in practice.
The third challenge in Koopman operator-based control is ensuring the stability of the Koopman operator itself since the real part of any continuous time eigenvalue of the Koopman operator of any physical system should not be positive [29, 31]. When the forward pass becomes unstable, the backward pass will be unstable as well. Lastly, curating training data for Koopman operator-based control often depends on ad-hoc selections of input signals and initial conditions, which can result in insufficient sampling of the high-dimensional phase space and, consequently, poor control performance.
In order to address the above issues, the goal of this paper is to develop a noisy-robust learning framework for Koopman operator with physics-informed observables. Specifically, the contributions of this work are:
  • approximating Koopman invariant subspace with physics-informed observables from Polyflow,
  • improving noise-robustness and long-term prediction accuracy via minimizing recurrent loss via roll-out,
  • stabilizing Koopman Operator by negative-definite parameterization,
  • iteratively learning Koopman-based model on augmented data with previous responses of the controlled nonlinear system.
The structure of the paper is as follows: Section II describes the proposed framework, including the use of Polyflow basis and the stabilized formulation. Section III presents the results of our experiments, demonstrating the efficacy of our approach. Section IV discusses the outcomes of applying iterative data augmentation on Koopman based control tasks. In Section V, we conduct an ablation study to evaluate the contributions of the various methods used in this paper. Section VI provides the conclusion of the paper.
II Proposed Framework
II-A Lifted Linear Predictor
Consider a nonlinear discrete-time dynamical system:
$x_{k+1}=f(x_{k},u_{k}),$ (1)
where $f(\cdot)$ represents a generally nonlinear transition mapping, and $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{m}$ denote the system’s state and input, respectively. The objective is to predict the trajectory of the system, given an initial condition $x_{0}$ at time $t_{0}$, and a sequence of inputs $\{u_{i}\}_{i=0}^{N_{p}}$, where $N_{p}$ is the prediction horizon length. To achieve this, we define the lifted linear predictor as a dynamical system:
$z_{0}$ $=\Phi(x_{0}),$ (2)
$z_{k+1}$ $=Az_{k}+Bu_{k},$
$\hat{x}_{k}$ $=Cz_{k},$
where $z\in\mathbb{R}^{N_{\phi}}$ is the lifted state, $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{N_{\phi}}$ is the lifting mapping, and $\hat{x}$ is the prediction of the original system’s state $x$. The lifting mapping $\Phi(\cdot)$ is defined as:
$\Phi(x)=[\phi_{1}(x),\phi_{2}(x),\ldots,\phi_{N_{\phi}}(x)]^{\top},$ (3)
where $\phi_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a scalar function referred to as an observable.
The predictor can also take the form of a bilinear system:
$z_{k+1}$ $=Az_{k}+(Bz_{k})u_{k},$ (4)
$\hat{x}_{k}$ $=Cz_{k},$
which can capture more complex dynamical behaviors. While our framework can be extended to accommodate the bilinear form, we avoid this for simplicity and to maintain readiness for application in linear control systems.
II-B Koopman Operator for Controlled System
To provide a rigorous justification for constructing a linear predictor via state-space lifting, we briefly outline the Koopman operator approach for analyzing controlled dynamical systems. We will skip the discussion for Koopman Operator for autonomous system.
We consider the following discrete-time nonlinear system with control:
$x^{+}$ $=f(x,u)$ (5)
where, $x\in\mathbb{R}^{n}$ is the state of the system and $u\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input. $\mathcal{U}$ is the space of the available control inputs. $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is the transition mapping. $x^{+}$ is the system state of the next time step.
We define the extended state comprising of the state $x$ of system and the control sequence $\mathbf{u}\in\ell(\mathcal{U})$, where $\ell(\mathcal{U})$ is the space of all sequences $(u_{i})_{i=0}^{\infty}$ with $(u_{i})_{i=0}^{\infty}\in\mathcal{U}$,
$\chi=\begin{bmatrix}x\\ \mathbf{u}\end{bmatrix},$
The dynamics of the extended state $\chi$ is described by
$\chi^{+}=F(\chi):=\begin{bmatrix}f(x,u(0))\\ \mathcal{S}\mathbf{u}\end{bmatrix},$ (6)
where $\mathcal{S}$ is the left shift operator, i.e., $(\mathcal{S}\mathbf{u})(i)=u(i+1)$, and $u(i)$ denotes the $i$th element of the sequence $\mathbf{u}$.
The Koopman operator $\mathcal{K}:\mathcal{H}\to\mathcal{H}$ associated to (6) is defined by
$(\mathcal{K}\psi)(\chi)=\psi(F(\chi))$
for any observable $\psi:\mathbb{R}^{n}\times\ell(\mathcal{U})\to\mathbb{R}$ belonging to some space of observables $\mathcal{H}$.
Despite Koopman operator $\mathcal{K}$ being a linear operator fully describing the time evolution of any observable function on the non-linear dynamical system, it is inherently infinite dimensional that makes learning Koopman operator challenging.
II-C Extended DMD for Control with Physics-Informed Observables
For practical purposes, we adapt to the Extended Dynamic Mode Decomposition (EDMD) [32] approach followed by Korda and Mezić [15]. The EDMD provides a straightforward method to approximate the Koopman operator from data, thereby obtaining the lifted linear predictor in (2). Since our goal is not to predict the control sequence, the observable function will depend only on the system states. Thus, the process begins by selecting $N_{\phi}$ observables $\phi_{1},\ldots,\phi_{N_{\phi}}$ and forming the predictor’s state $z$ from the measured state $x$ of the system (1) as follows:
$z=\Phi(x)=[\phi_{1}(x),\phi_{2}(x),\ldots,\phi_{N_{\phi}}(x)]^{\top}.$ (7)
The one-step prediction loss based EDMD algorithm assumes that a collection of data $(x_{j},u_{j},x_{j}^{+}),j=1,\ldots,K$ satisfying $x_{j}^{+}=f(x_{j},u_{j})$ is available and seeks matrices $A$ and $B$ minimizing
$\sum_{j=1}^{K}\|\Phi(x_{j}^{+})-A\Phi(x_{j})-Bu_{j}\|_{2}^{2},$ (8)
To select suitable basis functions $\Phi$ that span an invariant subspace, we are motivated by Polyflow basis functions defined in [26]. When the dynamics $f(x,u)$ is known, we can use a special set of basis functions $\Phi(x)=T_{N}(x)$, known as Polyflow basis functions of order $N\in\mathbb{Z}^{+}$ which is defined,
$T_{N}(x)$ $:=\begin{bmatrix}x\\ f(x,0)\\ f(f(x,0),0)\\ \vdots\\ f^{(N-1)}(x,0)\end{bmatrix}.$ (9)
Using these polyflow basis functions to lift the system state, we can derive the following system, which aligns with equation (2)
$z_{0}$ $=T_{N}(x_{0})$ (10)
$z_{k+1}$ $=Az_{k}+Bu_{k},$ (11)
$\hat{x}_{k}$ $=CT_{N}(x).$ (12)
Note that the matrix $C$ is typically known from the selection of $T_{N}(x)$.
Time delay embedding [33] and Polyflow offer distinct approaches to state lifting. Time delay embedding relies on observing and recording the physical system’s state over time to embed the state vector. In contrast, Polyflow directly utilizes the governing equations of the system, enabling an instantaneous embedding of the state without accessing the physical system itself. Our framework assumes the existence of a governing equation, but for systems with unknown dynamics, methods such as SINDy [34] or Neural ODE [35] can be employed to develop a working governing equation.
II-D Recurrent Roll-out Loss
To enhance long-term prediction accuracy, we implement a recurrent roll-out loss instead of using one-step prediction loss (8). For this, we consider a set of $M$ trajectories, each consisting of $R$ steps. Let $x_{i,j}$ denote the state at the $i$-th step of the $j$-th trajectory, where $i=0,1,\ldots,R$ and $j=1,2,\ldots,M$. The following optimization problem is then formulated:
$\underset{A,B}{\min}\sum_{j=1}^{M}\sum_{r=1}^{R}\|T_{N}(x_{r,j})-A^{r}T_{N}(x_{0,j})-\sum_{k=0}^{r-1}A^{r-k-1}Bu_{k,j}\|_{2}^{2}$ (13)
where, $A$ and $B$ are the matrices to be determined, and $u_{i}^{j}$ denotes the input at the $i$-th timestep of the $j$-th trajectory.
To determine an appropriate basis size $N$, we solve the optimization problem (13) iteratively, increasing $N$ until the approximation error is sufficiently small.
II-E Curriculum Learning
Curriculum learning is a training strategy where a model is first trained on simpler data points and then gradually introduced to more challenging ones. This approach helps the model to progressively learn and adapt to increasing complexity.
In our framework, we implement a form of curriculum learning by progressively increasing the roll-out length $R$ during the training process. Initially, the model is trained with shorter roll-out lengths, focusing on immediate predictions. As the training progresses and the model stabilizes, the roll-out length $R$ is increased geometrically after certain epochs. This gradual increase is expected to allow the model to learn from easier short-term predictions before tackling the more complex task of long-term prediction, thereby aligning with the curriculum learning paradigm. By doing so, the model is expected to build its capacity to handle extended sequences, improving overall prediction accuracy over long time horizons.
II-F Stabilized Formulation
To promote stability of the lifted system, we enforce stable eigenvalues of $A$ by employing the fact that a skew symmetric matrix with negative real valued diagonal elements have stable or neutral stable eigenvalues in continuous time dynamical system. For any $Q\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ matrix and $C\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ diagonal matrix with negative real diagonal elements, $\tilde{A}=Q-Q^{T}+C$ has eigenvalues in the negative real part of the complex plane. We compute the discrete-time $\hat{A}\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ by taking the matrix exponentiation of $\tilde{A}=Q-Q^{T}+C$ multiplied with the sampling timestep $\Delta t$.
$\hat{A}$ $=\exp\left(\tilde{A}\Delta t\right)$ (14)
A similarity transform of $\hat{A}$ via the coordinate transform matrix $P\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ is considered to find the Koopman Operator $A$.
$A$ $=P\hat{A}P^{-1}$ (15)
This formulation ensures the stability of the Koopman Operator throughout the training. We name this Dissipative Koopman Operator. In contrast, a Standard Koopman Operator Matrix can be any matrix $A$ without any stability constraint.
II-G Model Predictive Control
The application of lifted linear predictor within MPC is referred to as Koopman MPC (KMPC) in [15]. Here, we present the main formulation and omit the details, which can be found in the referenced literature. At each discrete time step, the MPC solves the following optimization problem over the prediction horizon $N_{p}$:
$\min_{u_{k},z_{k}}$ $J\left(\{z_{k}\}_{k=0}^{N_{p}},\{u_{k}\}_{k=0}^{N_{p}}\right),$ (16)
subject to $\quad z_{k+1}=Az_{k}+Bu_{k},\quad k=0,1,\ldots,N_{p}-1,$
$\quad e_{k}=r_{k}-Cz_{k},$
$\quad z_{\min}\leq z_{k}\leq z_{\max},$
$\quad u_{\min}\leq u_{k}\leq u_{\max},$
$\quad\text{parameters}\quad z_{0}=\Phi(x_{0}),$
$\quad r_{k}=\text{given},\quad k=0,1,\ldots,N_{p},$
where $u_{\min/\max}$ are the input bounds, $z_{\min/\max}$ are the bounds on the lifted states, and $r_{k}$ is the reference signal. Only the first control input $u_{0}$ from the optimal sequence is applied, and the optimization problem is resolved at the next time step. The cost function $J$ is given by:
$J(z,u)=e_{N_{p}}^{\top}Q_{N}e_{N_{p}}+\sum_{k=0}^{N_{p}-1}\left(e_{k}^{\top}Qe_{k}+u_{k}^{\top}Ru_{k}\right).$ (17)
II-H Iterative Learning
Models trained on trajectories that are generated by randomly sampling the phase space and control input can perform poorly when subjected to the optimal feedback control inputs. This is primarily because these trained models lack exposure to the desired scenarios during the training process, especially for relatively high-dimensional system. To address this issue, we propose an iterative data augmentation technique based on collecting response of the true system under control input that comes from the data-driven model. Such response together with control input is used to augment the original dataset to further refine and improve the model. Our rationale is that such experience from the true system can be used to better explore the phase space of nonlinear system than merely ad-hoc random sampling the phase space, which is subject to the curse of dimensionality.
Our proposed framework is named as STIRK (Stable Training with Iterative Roll-out Koopman Operator). A summary of the proposed framework is illustrated in the figure 1.
Figure 1: Overview of the roll-out EDMD Koopman Model. The process begins with trajectory data collected from a nonlinear dynamical system. The states are lifted through Polyflow basis functions to obtain a higher-dimensional representation. This representation is then used in the roll-out EDMD algorithm to derive a linear model. The resulting linear model is employed for Model Predictive Control (MPC), which is iteratively refined through iterative learning from closed loop trajectories.
III Numerical Experiments
In this section, we will evaluate our proposed framework by considering the following two tasks:
  • Prediction of the time evolution of Van der Pol oscillator given initial condition using the surrogate model learned with our framework;
  • Model predictive control of CartPole system given a surrogate model learned with our framework from noisy measurements.
III-A Prediction Performance
In order to evaluate the prediction performance of the proposed framework, we compare it with the following versions of DMD:
  1. 1.
    DMD [36],
  2. 2.
    eDMD  [32]with Polyflow basis function and Radial Basis function,
  3. 3.
    OptDMD [37] with Polyflow basis function, Radial Basis functions and Identity Observables,
  4. 4.
    TLSDMD [38] with Polyflow basis function, Radial Basis functions and Identity Observables,
  5. 5.
    FBDMD [39] with Polyflow basis function, Radial Basis functions and Identity Observables.
Van der Pol Oscillator
We consider the autonomous Van der Pol Oscillator described by the equations:
$\dot{x}_{1}$ $=-x_{2},$ (18)
$\dot{x}_{2}$ $=\mu\left(-1+x_{1}^{2}\right)x_{2}+x_{1},$ (19)
where $\mu$ is set to 1.0. The Van der Pol Oscillator is a nonlinear system characterized by a non-conservative force that can either dampen or amplify oscillations depending on the state. Since our framework relies on discrete-time dynamical system, the continuous-time version of Van der Pol system is discretized using a fourth-order Runge-Kutta method with a time step $\Delta t$ of 0.1. The training dataset is generated by running simulations for 100 steps, also with a time step of $\Delta t$ of 0.1. After the simulation, Gaussian noise is added to the states, with 10 noise levels logarithmically spaced from $10^{-3}$ to $10^{-1}$. For models trained on a single trajectory, the initial condition is sampled from a normal distribution with a mean of $[-1.0,-1.0]$ and a standard deviation of $[0.05,0.05]$. For models trained on multiple trajectories, 50 initial conditions are uniformly sampled from the range $[-1,1]$. To ensure robustness and generalizability, we create 10 sets of training data using different random seeds. Figure 2 illustrates one set of the training trajectories and test trajectories within the $x_{1},x_{2}$ plane.
Figure 2: Visualization of the Van der Pol system’s dynamical behavior, illustrating ground truth training trajectories, noisy data points (noise level 0.0599), and test trajectories in the $x_{1},x_{2}$ plane.
Prediction Performance Evaluation
To evaluate the prediction performance of the different algorithms, we use a normalized error metric that quantifies the discrepancy between the predicted trajectories and the ground truth trajectories. The evaluation is conducted as follows:
Let $\mathbf{x}(t)\in\mathbb{R}^{n}$ represent the true state of the system at time $t$ and $\hat{\mathbf{x}}(t)\in\mathbb{R}^{n}$ represent the predicted state at time $t$. The normalized prediction error over the trajectory is computed using the following formula:
$\text{Error}=\frac{\|\hat{\mathbf{X}}-\mathbf{X}\|_{F}}{\|\mathbf{X}\|_{F}},$ (20)
where $\|\cdot\|_{F}$ denotes the Frobenius norm, and $\mathbf{X}\in\mathbb{R}^{T\times n}$ and $\hat{\mathbf{X}}\in\mathbb{R}^{T\times n}$ are the matrices containing the true and predicted states over all time steps $T$, respectively,
$\mathbf{X}=\begin{bmatrix}\mathbf{x}(1)\\ \mathbf{x}(2)\\ \vdots\\ \mathbf{x}(T)\end{bmatrix},\quad\hat{\mathbf{X}}=\begin{bmatrix}\hat{\mathbf{x}}(1)\\ \hat{\mathbf{x}}(2)\\ \vdots\\ \hat{\mathbf{x}}(T)\end{bmatrix}.$ (21)
Training
The Van Der Pol system was initially trained using the Adam optimizer with a base learning rate of 0.01. A cyclic learning rate scheduler in triangular mode was employed, where the initial learning rate was halved each cycle. The maximum learning rate was set to 0.1, with each cycle stepping up for 500 epochs and stepping down for another 500 epochs. The Polyflow order was configured to 4. The training process spanned 15,000 epochs with a batch size of 1,000. The roll-out length was progressively increased, starting at 1 and doubling every 200 epochs, with a maximum roll-out length initially set to 90. In the final 1,000 epochs, the optimizer was switched to LBFGS with a learning rate of 0.01, applied on full batch data, to fine-tune the system.
Results for Single Trajectory
Figure 3(a) compares the reconstruction performance of various Dynamic Mode Decomposition (DMD)-based methods when trained on a single trajectory of the Van Der Pol system at a noise level of 0.0599. The legend lists the normalized error for each method. The Dissipative Koopman model with Polyflow basis functions (DISSIPATIVE-Polyflow) achieved the lowest error of $2.98\times 10^{-2}$, indicating superior accuracy. OptDMD, using both identity and Polyflow basis functions, performed similarly well, closely approaching the performance of the Dissipative Koopman model. However, since OptDMD also optimizes the initial condition, it shows a deviation from the actual initial condition. Other methods exhibited higher errors, indicating less accurate reconstructions under the given noise conditions. Figure 3(b) presents the mean squared error (MSE) trajectory over 100 timesteps for the Van Der Pol system at a noise level of 0.0599 for 10 different training sets. Each line represents the mean of the errors for a different DMD-based method, with shaded areas indicating the standard deviation of the error. The Dissipative Koopman model with Polyflow basis functions consistently maintains the lowest MSE across all timesteps, highlighting its robustness and accuracy under noisy conditions. The eDMD with Polyflow basis functions and the eDMD with RBF basis functions exhibit higher MSEs, particularly in the early timesteps, indicating less accurate predictions compared to the Dissipative model. The standard DMD with identity observables (DMD-Identity) also shows higher error compared to the Dissipative Koopman model but performs better than some of the other methods. The figures demonstrate that the Dissipative Koopman model with Polyflow basis functions outperforms other DMD-based methods in terms of prediction accuracy for the Van Der Pol system under noisy conditions. This highlights the effectiveness of incorporating dissipative properties and structured observables in enhancing the robustness and accuracy of the model.
(a) Single Trajectory Prediction for Van Der Pol System with Noise Level 0.0599. The plot compares the ground truth trajectory (black solid line) with the predictions from various DMD-based methods.
(b) Mean Squared Error (MSE) Trajectory over Timesteps for Van Der Pol System with Noise Level 0.0599. The plot shows the evolution of MSE over 100 timesteps for different DMD-based methods. Shaded areas represent mean $\pm$ standard deviation of the mean squared errors. The Dissipative Koopman model with Polyflow basis functions maintains the lowest error throughout the timesteps, demonstrating robust performance under noisy conditions.
Figure 3: Single Trajectory Prediction Performance For Van Der Pol System
Results for Multiple Trajectories
Figure 4 compares the prediction errors for multiple trajectories of the Van Der Pol (VDP) system across different noise levels. Panel (a) shows the training error, while panel (b) shows the test error. The Dissipative Koopman model with Polyflow basis functions consistently shows lower errors across various noise levels, demonstrating its robustness and accuracy in handling multiple trajectories.
(a) Training Error
(b) Test Error
Figure 4: Multiple Trajectory Error Comparison for VDP system
TABLE I: Summary of Prediction Performance for Van Der Pol System across different strategies
Strategy DMD FBDMD TLSDMD eDMD DISSIPATIVE (Ours)
Lifting Function Identity Identity Polyflow RBF Identity Polyflow RBF Polyflow RBF Polyflow
Error Type Noise
Mean Normalized Training Error 0.0010 0.1198 0.1211 0.0697 0.1079 0.1210 NaN NaN 0.0547 0.0898 0.0417
0.0599 0.5660 0.1638 NaN 0.1468 0.1249 NaN NaN 0.5225 0.4500 0.0547
0.1000 0.7711 0.2647 NaN 0.2486 0.1348 NaN NaN 0.7149 0.6420 0.0733
Mean Normalized Test Error 0.0010 0.1315 0.1326 0.1028 0.1268 0.1325 NaN NaN 0.0796 0.1085 0.0590
0.0599 0.5704 0.1691 NaN 0.1671 0.1370 NaN NaN 0.5370 0.4688 0.0694
0.1000 0.7715 0.2634 NaN 0.2718 0.1468 NaN NaN 0.7235 0.6552 0.0888
III-B Control Performance
CartPole System
Let us consider the CartPole System,
$\dot{x}_{1}$ $=x_{3},$ (22)
$\dot{x}_{2}$ $=x_{4},$ (23)
$\dot{x}_{3}$ $=\frac{u+m_{p}\sin x_{2}\left(lx_{4}^{2}-g\cos x_{2}\right)}{m_{c}+m_{p}\sin^{2}x_{2}},$ (24)
$\dot{x}_{4}$ $=\frac{u\cos x_{2}+m_{p}lx_{4}^{2}\cos x_{2}\sin x_{2}-(m_{c}+m_{p})g\sin x_{2}}{l(m_{c}+m_{p}\sin^{2}x_{2})}.$ (25)
The CartPole system consists of a cart that moves along a track while balancing a pole attached to it via a pivot. The dynamics of the system are described by the equations above, where $x_{1}$ represents the position of the cart, $x_{2}$ represents the angle of the pole, $x_{3}$ represents the velocity of the cart, $x_{4}$ represents the angular velocity of the pole, and $u$ is the control input force applied to the cart. The parameters of the system include the mass of the cart ($m_{c}$), the mass of the pole ($m_{p}$), the length of the pole ($l$), and the acceleration due to gravity ($g$). For our framework, the continuous-time CartPole system is discretized using a fourth-order Runge-Kutta method with a time step $\Delta t$ of 0.05. The training dataset is generated by running simulations for 200 steps and contains 50 different trajectories. After the simulation, Gaussian noise is added to the states with noise level of 0.1. The initial conditions for the CartPole system are randomly sampled with the position ($x_{1}$) from $[-1,1]$, the angle ($x_{2}$) from $[-\pi/2,\pi/2]$, the velocity ($x_{3}$) from $[-0.1,0.1]$, and the angular velocity ($x_{4}$) from $[-0.1,0.1]$. Control inputs are generated using exponentially decaying sine waves of randomly sampled frequencies, amplitudes and phase angles.
Training
The CartPole system was trained using the Adam optimizer with a learning rate of 0.001. The Polyflow order was set to 4. The training dataset consisted of 45 trajectories, while 5 trajectories were reserved for validation. The training process spanned 5,000 epochs with a batch size of 2,000. The maximum roll-out length was set to 32 and the roll-out length was increased every 300 epochs by a factor of 2.
MPC Configuration
The trained Koopman model is used as a surrogate model for model predictive control, following the approach in [15]. The prediction horizon, $N_{P}$, is set to 20 steps. Control input constraints are defined as $u_{\text{min}}=-20$ and $u_{\text{max}}=20$, with no constraints on the system states. For the cost function described in (17), $Q_{N_{P}}$ and $Q_{N}$ are set to ones on the first $n$ diagonal elements, with zeros elsewhere. The parameter $R$ is set to 0.
Results
Figure 5 compares the Model Predictive Control (MPC) costs for the CartPole system using three different models under a noise level of 0.100. The models compared are EDMD with RBF basis functions (eDMD-RBF), the Dissipative Koopman model with Polyflow basis functions (Dis-PF), and the Standard Koopman model with Polyflow basis functions (Std-PF). The violin plots illustrate the distribution of MPC costs across various initial states. The results indicate that both the Dissipative Koopman model with Polyflow basis functions (Dis-PF) and the Standard Koopman model with Polyflow basis functions (Std-PF) achieve relatively similar MPC costs. Both models show stable control performance across different initial states, demonstrating robustness in control. In contrast, the EDMD with RBF basis functions (eDMD-RBF) exhibits higher variability in MPC costs and generally incurs higher average MPC costs compared to the Dissipative and Standard Koopman models. This suggests that while the EDMD-RBF model provides some level of stability, it does so at the expense of higher control costs and less consistency. Overall, the comparison reveals that both the Dissipative and Standard Koopman models with Polyflow basis functions offer superior and more stable control performance compared to the EDMD-RBF model under the noise level of 0.1000. This highlights the effectiveness of the Polyflow basis functions in achieving robust control in MPC applications.
Figure 5: Violin plots showing the distribution of Model Predictive Control (MPC) costs for different models (eDMD-RBF, Dis-PF, Std-PF) under various initial states.
IV Iterative Data Augmentation
To more effectively explore the phase space of a nonlinear system, we introduce an iterative data augmentation technique. Instead of relying solely on ad-hoc random sampling, we augment the training data with failure trajectories from closed-loop control and then retrain the Koopman operator. This approach addresses challenges highlighted by [40], where the authors emphasize the need to identify inputs that closely approximate the final closed-loop input and to use randomized signals that resemble it. However, finding suitable randomized control signals can be challenging and may not be adequate for control tasks. Therefore, we propose a rather automated approach: augmenting the training data with control signals from previous failed control tasks in a closed-loop setting. This approach involves collecting the system’s response to control inputs from data-driven based model predictive controller. The responses, along with the corresponding control inputs, are used to enhance the original dataset, enabling further refinement and improvement of the model.
Figure 6: Trajectory visualization of the Cartpole system over three learning iterations. The starting positions are indicated in light grey, while the final positions are highlighted in red.
Figure 6 visualizes the trajectory of the CartPole system across three different scenarios: no data augmentation, 1x data augmentation, and 2x data augmentation. In the first scenario, where no data augmentation was applied, the Koopman model fails to stabilize the CartPole after 200 steps. However, when the failed trajectory, along with other failed trajectories from different initial conditions, is added to the training data, the Koopman model demonstrates significantly improved control outcomes. The system’s performance notably improves with data augmentation, transitioning from oscillating to stable trajectories. This enhancement is also reflected in the reduction of the Model Predictive Control (MPC) cost, with costs decreasing from 135.6 (no augmentation) to 74.7 (1x augmentation) and 67.8 (2x augmentation).
Figure 7: Violin plots depicting the distribution of Model Predictive Control (MPC) costs for Cartpole problem for various tasks and iterative models under different initial states for noise level 0.1000.
Figure 7 presents violin plots illustrating the distribution of Model Predictive Control (MPC) costs for the CartPole system across two different tasks—seen and unseen—and iterative models under a noise level of 0.1000. The “seen” task refers to scenarios where input bounds are set between $[-20,20]$, while the “unseen” task refers to tighter input bounds of $[-10,10]$. Data augmentation was performed using closed loop trajectories from the seen task. The violin plots display MPC costs for four different initial states and two levels of data augmentation, comparing the costs for both control tasks. Each subplot emphasizes the performance gains achieved through iterative learning. The results indicate that both seen and unseen control tasks benefit significantly from iterative learning, with successive iterations leading to lower and more stable MPC costs. This demonstrates the effectiveness of iterative learning in refining control strategies and enhancing the overall performance of the system.
V Ablation Studies
In this section, we will analyze the individual effect of the following components within our framework on the overall model performance with training data at different noise levels:
  • Stabilized formulation of Koopman operator in Section II-F,
  • Learning rate scheduler,
  • Optimizer switching,
  • Progressive roll-out.
V-A Dissipative Koopman Model versus Standard Koopman Model
Figure 9(a) compares the performance of the Dissipative Koopman Model and the Standard Koopman Model across different noise levels for multiple trajectories. The Dissipative Koopman Model consistently shows lower mean errors and mean test errors, underscoring its superior prediction accuracy and robustness. However, the training time for the Dissipative Koopman Model is considerably longer than that of the Standard Koopman Model, due to the increased number of learning parameters and the computationally intensive operations, such as matrix exponentiation and pseudo-inverse computation. Despite its relatively short training time, a significant drawback of the Standard Koopman Model is its sensitivity to the initial conditions of the Koopman matrix. If these initial conditions are not carefully tuned, the model can produce unstable eigenvalues, leading to poor performance. Figure 8 illustrates this issue by showing the discrete-time eigenvalues of the standard and dissipative Koopman operators for the Van Der Pol Oscillator System at a noise level of 0.0599, using a different weight initialization than in previous cases. For fair comparison, both standard and dissipative Koopman models are initialized with stable Koopman matrices. However, after 1,000 training epochs, some of the discrete-time eigenvalues in the Standard Koopman matrix move outside the unit circle, leading to a loss of stability while dissipative Koopman model never loses stability during training.
Figure 8: Eigenvalues of the Standard and Dissipative Koopman operators for the Van Der Pol system with a noise level of 0.0599 after 1,000 training epochs
V-B Learning Rate Scheduler
Figure 9(b) compares the performance of the model using constant learning rate (LR) scheduling and cyclic learning rate scheduling under different noise levels. The cyclic learning rate scheduling generally results in slightly lower mean errors and mean test errors compared to the constant learning rate. The time taken for training with cyclic learning rate scheduling is slightly higher due to the oscillation in learning rates.
V-C Optimizer Switching
Figure 9(c) examines the effect of switching optimizers on the model’s performance. Different combinations of Adam and LBFGS optimizers were used, and their impact on mean errors, mean test errors, and training times was analyzed. Switching optimizers can lead to lower mean errors and test errors in some configurations, suggesting that optimizer switching can enhance training effectiveness. The combination of Adam and LBFGS every 500 epochs generally shows better performance with lower errors but increases the training time significantly. The mean error and test error for configurations where optimizers are switched show improved stability and reduced overfitting.
V-D Progressive roll-out Length Vs. Constant roll-out Length
Figure 9(d) compares the performance of the model using progressive roll-out length scheduling and constant roll-out length scheduling under different noise levels. The progressive roll-out scheduling shows lower mean errors and mean test errors compared to the constant roll-out length, indicating better learning and generalization. The time taken for training with progressive roll-out scheduling is higher due to the incremental increase in roll-out length.
(a) Effect of stabilized formulation of the Koopman operator.
(b) Effect of learning rate scheduler.
(c) Effect of optimizer switching.
(d) Effect of progressive roll-out.
Figure 9: Ablation study on the Van Der Pol Oscillator system. First column: training error. Second column: testing error. Third column: computational wall time.
VI Conclusions
In this work, we proposed a novel data-driven learning framework for the Koopman operator of nonlinear dynamical systems. Unlike existing frameworks that rely on ad-hoc observables or complex neural networks to approximate the Koopman invariant subspace, our approach utilizes observables chosen from known system dynamics with closure through the concept of Polyflow. This physics-informed basis significantly enhances the performance and robustness of the resulting models compared to ad-hoc choices. To ensure noise robustness and long-term stability of the lifted system, we designed a stable parameterization of the Koopman operator and employed a roll-out recurrent loss. Our stabilized parameterization guarantees a stable system, which can be equally accurate or, in some cases, even more accurate than the standard Koopman operator. Additionally, we addressed the curse of dimensionality in sampling from high-dimensional phase space by iteratively augmenting data from closed-loop trajectories. This approach not only improved performance in seen closed-loop scenarios but also enhanced the model’s ability to handle unseen control objectives effectively. We conducted an ablation study to demonstrate the effectiveness of several key components: switching to the LBFGS optimizer for fine-tuning, employing cyclic learning rate scheduling, utilizing progressive roll-out lengths for enhanced training outcomes, and incorporating a stabilized Koopman operator to ensure guaranteed system stability. Our numerical results on the Van der Pol Oscillator and the CartPole system demonstrated the superior performance of the proposed framework in prediction and model predictive control (MPC) of nonlinear systems, compared to other popular DMD-based methods Overall, our framework offers a noise-robust approach for learning stable Koopman operators, providing a solid foundation for future research and applications in nonlinear dynamical system control.
Acknowledgement
The authors are thankful for the provision of computational resources from the Center for Computational Innovations (CCI) at Rensselaer Polytechnic Institute (RPI) during the early stages of this research. Numerical experiments are performed using the computational resources granted by NSF-ACCESS for the project PHY240112 and that of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility using NERSC award NERSC DDR-ERCAP0030714.